When the Theban legion refused to obey the orders, it was decreed that the legions should be decimated-every tenth man killed. If 590 men were killed, how many men did the Theban legion have after it was decimated?

Answer by Theo(13342)

if every tenth man was killed and 590 were killed, then the formula for how many there were originally would be:
1/10 * x = 590, x representing the original number of thebans.
solve for x to get:
x = 10 * 590 = 5900
subtract the 590 that were killed to get a remaining number of thebans left equal to 5900 - 590 = 5310.
that should be your solution.
here's a link to some information on the history of the theban legion.
https://catholicsaints.info/martyrs-of-the-theban-legion/
it appears that killing 1 out of every 10 of them was only the beginning.
they were all eventually slaughtered.